OpenStudy (anonymous):

What are the first five terms of the sequence given by the formula an = 5n + 1? 1, 5, 9, 13, 17 6, 11, 16, 21, 26 5, 10, 15, 20, 25 5, 6, 7, 8, 9

OpenStudy (raden):

an = 5n + 1 for x=1 then a1 = 5(1) + 1 = 5 + 1 = 6 for x=2 then a2 = 5(2) + 1 = 10 + 1 = 11 so on ....